# Decentralized Parametric Stability Certificates for Grid-Forming Converter Control

### v5

@@ III. Passivity Checks & Theorem 1
$\begin{split}a_{0,i}&=\omega_{0}^{2}(\rho^{4}+2\rho^{2}-6\rho\alpha_{\mathrm{p},i}+1)\\ a_{1,i}&=\omega_{0}(\tilde{\tau}_{\mathrm{p},i}\rho^{4}+2\rho^{3}+2\rho^{2}\tilde{\tau}_{\mathrm{p},i}\\ &-2\rho^{2}\alpha_{\mathrm{p},i}+2\rho-5\alpha_{\mathrm{p},i}+\tilde{\tau}_{\mathrm{p},i})\\ a_{2,i}&=(1+\rho^{2})(2\tilde{\tau}_{\mathrm{p},i}\rho+1)\\ a_{3,i}&=\tau_{\mathrm{p},i}(1+\rho^{2})\end{split}\begin{split}b_{0,i}&=\omega_{0}^{2}(1+\rho^{2})^{2}\\ b_{1,i}&=2\rho\omega_{0}(1+\rho^{2})\\ b_{2,i}&=1+\rho^{2}.\end{split}$ (39)
By checking the Hurwitz criterion for a 3rd order polynomial, we can conclude that $a_{3,i}>0$ and $a_{2,i}>0$ are always satisfied, while $a_{1,i}>0$ is ensured by condition 24a, $a_{0,i}>0$ by condition 24b, and $a_{2,i}a_{1,i}>a_{0,i}a_{3,i}$ by condition 24c.
(ii) Positive definiteness: To ensure $\mathcal{D}_{\mathrm{p},i}^{\prime}(\mathrm{j}\omega)+\mathcal{D}^{\prime\star}_{\mathrm{p},i}(\mathrm{j}\omega)\succ 0$, we require the additional condition 24d.
Following the same reasoning, we derive the decentralized stability conditions [^3] The requirement $\tau_{\mathrm{q},i}\neq 0$ in 25 ensures properness of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$. for the reactive power-voltage droop control presented in 25 by checking strict passivity of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
Finally, since $\mathcal{N}^{\prime}(s)$ is passive and $\mathcal{D}^{\prime}(s)$ is strictly passive, we can apply Theorem 1 and conclude $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
@@ IV. Final Value Theorem & Lemma 2
The internal feedback stability of the original system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ follows directly from the internal feedback stability of the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, as stated in the following lemma.
@@
Consider the device and network dynamics $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ in 28, as well as $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 23 with the network model 17. Then, internal feedback stability of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ implies internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$.
@@
We first conclude that $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ implies $(I+\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}=(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\in\mathcal{RH}_{\infty}^{2n\times 2n}$. Further, since
@@
$(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$, and we can thus conclude stability of $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$ if the step response of the voltage derivatives converges to zero, i.e.,
@@
To show that 41 holds, we first rewrite the transfer matrix $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=(I+\mathcal{D}{N}\tfrac{1}{s})^{-1}\mathcal{D}$, and then use the row permutation matrix $\mathcal{P}\in\mathbb{R}^{2n\times 2n}$ with elements
@@
$\mathcal{D}^{\mathcal{P}}(s)=\begin{bmatrix}\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}(s)&0_{n\times n}\\ 0_{n\times n}&\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}(s)\end{bmatrix},\,\,N^{\mathcal{P}}(s)=\begin{bmatrix}N_{1}^{\mathcal{P}}(s)&N_{2}^{\mathcal{P}}(s)\\ -N_{2}^{\mathcal{P}}(s)&N_{3}^{\mathcal{P}}(s)\end{bmatrix}$ (43)
@@
Recalling that $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}$ is stable, and thus its permuted version, we can apply the Final Value Theorem (FVT):
$\begin{split}&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(I+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s)\tfrac{1}{s})^{-1}\mathcal{D}^{\mathcal{P}}(s)\\ =\,&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(sI+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s))^{-1}\mathcal{D}^{\mathcal{P}}(s)s\\ =\,&\,\,\underset{s\rightarrow 0}{\text{lim}}\,\left[\begin{array}[]{c}(sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s){N}^{\mathcal{P}}_{1}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s)s\\ 0_{n\times n}\end{array}\right.\\ \quad&\hskip 71.13188pt\left.\begin{array}[]{c}0_{n\times n}\\ (sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s){N}^{\mathcal{P}}_{3}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s)s\end{array}\right]\\ =\,&\begin{bmatrix}\star_{n\times n}&0_{n\times n}\\ 0_{n\times n}&0_{n\times n}\end{bmatrix},\end{split}$ (46)
where for the second equality we have used $N_{2}^{\mathcal{P}}(0)=0_{n\times n}$. The last equality follows from the fact that $N_{1}^{\mathcal{P}}(0)$ is a Laplacian matrix with zero eigenvalue [22], while $N_{3}^{\mathcal{P}}(0)$ is a regular matrix which does not have a zero eigenvalue. Given 46, 41 follows, and thus $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
@@
(a) Stability conditions in 24.
(b) Stability conditions in 25.
Figure 9: 2D plot of the stability conditions for the three-node system with $\rho=0.05$. The stars indicate the feasibility of the GFM 3 control parameters.
@@ 5.1 Case Study I: Linearized & reduced models
We numerically validate Theorem 2 and Corollary 1 with MATLAB/Simulink using linearized small-signal models of a three-node system with three GFM VSCs. We consider a network with uniform resistance-inductance ratio $\rho=0.05$, a maximum steady-state voltage magnitude $|v|_{\mathrm{max}}=1.1$ pu, and identical self susceptances $\sum_{i\neq j}^{3}b_{ij}=5$ pu of all VSC nodes $i\in\{1,2,3\}$. The simulations follow the ideal (i.e., linearized and reduced) block diagram dynamics in Fig. 4. The nodes GFM 1 and GFM 2 employ fixed controllers that always satisfy the conditions in 24 and 25, while for node GFM 3 we are exploring varying control parameters.
@@
We begin by modeling the network dynamics $\mathcal{N}_{0}(s)$ using Network-Simplification Level 2 in 19, without imposing additional conditions on the controller of GFM 3 (see “no cond.” in Fig. 9). A small-signal load disturbance at node 1 reveals that the closed-loop system remains stable (Fig. 10), confirming the validity of Corollary 1. Likewise, when modeling $\mathcal{N}_{0}(s)$ with Network-Simplification Level 1 in 18 and ensuring that the controller of GFM 3 satisfies the stability conditions in 26 (see “cond. L1” in Fig. 9), we again observe closed-loop stability (Fig. 11). This further confirms Corollary 1. Our main result, Theorem 2, is validated in Fig. 12, where we model $\mathcal{N}_{0}(s)$ dynamically as in 17, and equip GFM 3 with a controller that meets the stability conditions in 24 and 25 (see “cond. DYN” in Fig. 9). Stability is immediately evident. In contrast, using a GFM 3 controller that satisfies only 26 (“cond. L1”) or no conditions at all (“no cond.”) leads to instability in the dynamic network model. This underscores the importance of accurate network modeling in control design, which is overlooked in the overly optimistic stability assessment of [7, 9, 10].

### v6

@@ III. Passivity Checks & Theorem 1
$\begin{split}\hskip-2.84526pta_{0,i}&\hskip-1.42262pt=\hskip-1.42262pt\omega_{0}^{2}(1\hskip-1.42262pt+\hskip-1.42262pt\rho^{2})-2\rho\omega_{0}d_{\mathrm{p},i}\gamma_{3,i}^{\mathrm{p}}\\ \hskip-2.84526pta_{1,i}&\hskip-1.42262pt=2\rho\omega_{0}+\tau_{\mathrm{p,i}}\omega_{0}^{2}(1\hskip-1.42262pt+\hskip-1.42262pt\rho^{2})\hskip-1.42262pt\\ &\hskip-1.42262pt-d_{\mathrm{p},i}\omega_{0}^{2}\gamma_{1,i}^{\mathrm{p}}-d_{\mathrm{p},i}\gamma_{3,i}^{\mathrm{p}}\\ \hskip-2.84526pta_{2,i}&\hskip-1.42262pt=\hskip-1.42262pt2\omega_{0}{\tau}_{\mathrm{p},i}\rho\hskip-1.42262pt+\hskip-1.42262pt1\\ \hskip-2.84526pta_{3,i}&\hskip-1.42262pt=\hskip-1.42262pt\tau_{\mathrm{p},i}\end{split}\begin{split}b_{0,i}&\hskip-1.42262pt=\hskip-1.42262pt\omega_{0}^{2}(1\hskip-1.42262pt+\hskip-1.42262pt\rho^{2})\hskip-2.84526pt\\ b_{1,i}&\hskip-1.42262pt=\hskip-1.42262pt2\rho\omega_{0}\hskip-2.84526pt\\ b_{2,i}&\hskip-1.42262pt=\hskip-1.42262pt1.\end{split}$ (40)
(i) Poles: To show that the poles of all elements of $\mathcal{D}_{\mathrm{p},i}^{\prime}(s)$ are in $\text{Re}(s)<0$, we check the Hurwitz criterion for a 3rd order polynomial, and conclude that $a_{3,i}>0$ and $a_{2,i}>0$ are always satisfied, while $a_{1,i}>0$ is ensured by 25a, $a_{0,i}>0$ by 25b, and $a_{2,i}a_{1,i}>a_{0,i}a_{3,i}$ by 25c.
(ii) Positive definiteness: To ensure $\mathcal{D}_{\mathrm{p},i}^{\prime}(\mathrm{j}\omega)+\mathcal{D}^{\prime\star}_{\mathrm{p},i}(\mathrm{j}\omega)\succ 0,\,\forall\omega\in(-\infty,\infty)$, we require the additional condition 25d.
Following the same reasoning, we derive the decentralized stability conditions for the reactive power-voltage droop control in 26 by checking strict passivity of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$. To do so, we select $\tfrac{d_{\mathrm{q},i}}{|v|_{0,i}}=\tfrac{1}{\tilde{\gamma}_{3,i}^{\mathrm{q}}}$ to cancel the zero at the origin, and ensure strict passivity of the minimal realization of $\mathcal{D}_{\mathrm{q},i}^{\prime}(s)$.
Finally, since $\mathcal{N}^{\prime}(s)$ is passive and $\mathcal{D}^{\prime}(s)$ is strictly passive, and additionally $\bar{\sigma}(\mathcal{N}^{\prime}(\mathrm{j}\infty))\bar{\sigma}(\mathcal{D}^{\prime}(\mathrm{j}\infty))<1$ (because $\mathcal{N}^{\prime}(\mathrm{j}\infty)=0_{2n\times 2n}$), we can apply Theorem 1 and conclude $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
@@ IV. Final Value Theorem & Lemma 2
The internal feedback stability of the original system $\mathcal{D}_{0}\#\mathcal{N}_{0}$ follows directly from the internal feedback stability of the minimal realization of all four closed-loop transfer functions of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, as stated in the following lemma.
@@
Consider the device and network dynamics $\mathcal{D}^{\prime}(s)$ and $\mathcal{N}^{\prime}(s)$ in 29, as well as $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$ in 24 with the network model 18. Then, internal feedback stability of the minimal realization of all four closed-loop transfer functions of the loop-shifted system $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$ implies internal feedback stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}$.
@@
We first conclude that $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ implies $(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}\hskip-1.42262pt=\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$, where the latter closed-loop transfer functions are minimal realizations. Further, since
@@
$(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt=\hskip-1.42262pt\text{diag}(1,\tfrac{s}{|v|_{0,1}},\dots,1,\tfrac{s}{|v|_{0,n}})(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$, and we can thus conclude stability of $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}$ if the step response of the voltage derivatives converges to zero, i.e.,
@@
To show that 42 holds, we first rewrite the transfer matrix $(I+\mathcal{D}\mathcal{N})^{-1}\mathcal{D}=(I+\mathcal{D}{N}\tfrac{1}{s})^{-1}\mathcal{D}$, and then use the row permutation matrix $\mathcal{P}\in\mathbb{R}^{2n\times 2n}$ with elements
@@
$\hskip-2.84526pt\mathcal{D}^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt=\hskip-2.84526pt\begin{bmatrix}\mathcal{D}_{\mathrm{p}}^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt&0_{n\times n}\\ 0_{n\times n}&\hskip-1.42262pt\mathcal{D}_{\mathrm{q}}^{\mathcal{P}}\hskip-1.42262pt(s)\end{bmatrix}\hskip-2.84526pt,\,\,N^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt=\hskip-2.84526pt\begin{bmatrix}N_{1}^{\mathcal{P}}(s)\hskip-1.42262pt&\hskip-1.42262ptN_{2}^{\mathcal{P}}\hskip-1.42262pt(s)\\ -N_{2}^{\mathcal{P}}\hskip-1.42262pt(s)\hskip-1.42262pt&\hskip-1.42262ptN_{3}^{\mathcal{P}}(s)\end{bmatrix}\hskip-2.84526pt$ (44)
@@
Recalling that $(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}$ is stable, and thus its permuted version, we can apply the Final Value Theorem (FVT):
$\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(I+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s)\tfrac{1}{s})^{-1}\mathcal{D}^{\mathcal{P}}(s)$ (47a)
$=\,$ $\,\,\underset{s\rightarrow 0}{\text{lim}}\,\,(sI+\mathcal{D}^{\mathcal{P}}(s){N}^{\mathcal{P}}(s))^{-1}\mathcal{D}^{\mathcal{P}}(s)s$ (47b)
$=\,$ $\,\,\underset{s\rightarrow 0}{\text{lim}}\,\left[\begin{array}[]{c}\hskip-4.26773pt(sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s){N}^{\mathcal{P}}_{1}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{p}}(s)s\\ \hskip-4.26773pt0_{n\times n}\end{array}\right$. (47e)
$\hskip 71.13188pt\left.\begin{array}[]{c}0_{n\times n}\hskip-4.26773pt\\ (sI+\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s){N}^{\mathcal{P}}_{3}(s))^{-1}\mathcal{D}^{\mathcal{P}}_{\mathrm{q}}(s)s\hskip-4.26773pt\end{array}\right]\hskip-2.84544pt$ (47h)
$=\,$ $\begin{bmatrix}\star_{n\times n}&0_{n\times n}\\ 0_{n\times n}&0_{n\times n}\end{bmatrix}$, (47i)
where for 47h we have used $N_{2}^{\mathcal{P}}(0)=0_{n\times n}$. The last equality in 47i follows from the fact that $N_{1}^{\mathcal{P}}(0)$ is a Laplacian matrix with zero eigenvalue [26], while $N_{3}^{\mathcal{P}}(0)$ is a regular matrix which does not have a zero eigenvalue. Given 47, 42 follows, and thus $(I+\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\in\mathcal{RH}_{\infty}^{2n\times 2n}$.
@@
Remark 6.
Recall again that we concluded stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$ by first proving stability of the minimal realization of all four closed-loop transfer functions of $\mathcal{D}^{\prime}\#\mathcal{N}^{\prime}$, which implies stability of $(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}^{\prime}\mathcal{N}^{\prime})^{-1}\mathcal{D}^{\prime}\hskip-1.42262pt=\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$, where the latter closed-loop transfer functions are minimal realizations. Notice that there are RHP pole-zero cancellations between $\mathcal{D}(s)$ and $\mathcal{N}(s)$, i.e., we can NOT conclude stability of $\mathcal{D}\#\mathcal{N}$. However, from the stability of $\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}\mathcal{N})^{-1}\mathcal{D}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$, we can conclude stability of $\hskip-1.42262pt(I\hskip-1.42262pt+\hskip-1.42262pt\mathcal{D}_{0}\mathcal{N}_{0})^{-1}\mathcal{D}_{0}\hskip-1.42262pt\in\hskip-1.42262pt\mathcal{RH}_{\infty}^{2n\times 2n}$ (cf. proof of Lemma 4), and, given there are no RHP pole-zero cancellations between $\mathcal{D}_{0}(s)$ and $\mathcal{N}_{0}(s)$, we conclude stability of $\mathcal{D}_{0}\#\mathcal{N}_{0}\in\mathcal{RH}_{\infty}^{4n\times 4n}$.
@@ 5 Numerical Case Studies
To validate our theoretical stability guarantees, we conduct numerical case studies in MATLAB/Simulink using a simple three-bus system with three GFM VSCs. This deliberately minimal setup enables a clear interpretation and verification of the theoretical results without the added complexity of large-scale systems. Specifically, it allows a comparison between (i) an analytical model derived from ideal block-diagram representations of the network and device dynamics, used to obtain the stability certificates (Case Study I), and (ii) a nonlinear electromagnetic transient (EMT) simulation with more realistic models (Case Study II). Since the proposed stability certificates are fully decentralized, they are independent of network size and the number of units. Hence, while we focus on a small system for clarity, the results can be readily extended to larger systems, a direction we leave for future work.
@@ 5.1 Case Study I: Linearized & reduced models
We numerically validate Theorem 2 and Corollary 1 using linearized small-signal models of a three-bus system with three GFM VSCs, as shown in Fig. 10 with parameters in Table 2(b). We consider a network with uniform resistance-inductance ratio $\rho=0.05$, a maximum steady-state voltage magnitude $|v|_{\mathrm{max}}=1.1$ pu, and identical self susceptances $\sum_{i\neq j}^{3}b_{ij}=5$ pu of all VSC nodes $i\in\{1,2,3\}$. The simulations follow the ideal (i.e., linearized and reduced) block-diagram dynamics in Fig. 4. GFM 1 and GFM 2 employ fixed controllers that always satisfy the conditions in 25 and 26, while for GFM 3 we are exploring varying control parameters.
[…]
Figure 10: Schematic one-line diagram of the three-bus power system grid topology used in numerical case studies.
Table 2: Network and GFM parameters
(a) Electrical parameters & set points.
Parameter Symbol Value
Base power, voltage, frequency $S_{\mathrm{b}}$, $V_{\mathrm{b}}$, $f_{\mathrm{b}}$ 100 MW, 230 kV, 50 Hz
RLRL line components $r_{ij}$, $l_{ij}$ 0.02, 0.4 pu
Shunt capacitors $c_{ij}$ 0.1 pu
Constant impedance loads $p_{\mathrm{load},i}$ 0.1 pu
Converter power ratings $S_{1}$, $S_{2}$, $S_{3}$ 100, 100, 200 MW
Converter $RL$ filter components $r_{\mathrm{f},i}$, $l_{\mathrm{f},i}$ 0.01, 0.1 pu
Converter steady-state voltages $|v|_{0,1}$$|v|_{0,2}$$|v|_{0,3}$ 1, 0.9, 1.1 pu
Converter power set points $p_{0,1}$, $p_{0,2}$, $p_{0,3}$ 0.1025, 0.1026, 0.0526 pu
Converter power set points $q_{0,1}$, $q_{0,2}$, $q_{0,3}$ 0, -0.635, 0.3495 pu
(b) GFM converter parameters.
Parameter Symbol Value
Droop gains GFM 1 (fixed) $d_{\mathrm{p,1}}$, $d_{\mathrm{q,1}}$ 0.003 pu,0.01 pu
Time constants GFM 1 (fixed) $\tau_{\mathrm{p,1}}$, $\tau_{\mathrm{q,1}}$ 0.1 s, 0.1 s
Droop gains GFM 2 (fixed) $d_{\mathrm{p,2}}$, $d_{\mathrm{q,2}}$ 0.003 pu,0.01 pu
Time constants GFM 2 (fixed) $\tau_{\mathrm{p,2}}$, $\tau_{\mathrm{q,2}}$ 0.1 s, 0.1 s
Droop gains GFM 3 (no cond.) $d_{\mathrm{p,3}}$, $d_{\mathrm{q,3}}$ 0.25 pu, 0.25 pu
Time constants GFM 3 (no cond.) $\tau_{\mathrm{p,3}}$, $\tau_{\mathrm{q,3}}$ 0 s, 0 s
Droop gains GFM 3 (cond. L1) $d_{\mathrm{p,3}}$, $d_{\mathrm{q,3}}$ 0.2 pu, 0.2 pu
Time constants GFM 3 (cond. L1) $\tau_{\mathrm{p,3}}$, $\tau_{\mathrm{q,3}}$ 0.01 s, 0.01 s
Droop gains GFM 3 (cond. DYN) $d_{\mathrm{p,3}}$, $d_{\mathrm{q,3}}$ 0.006 pu, 0.02 pu
Time constants GFM 3 (cond. DYN) $\tau_{\mathrm{p,3}}$, $\tau_{\mathrm{q,3}}$ 0.1 s, 0.1 s
@@
We begin by modeling the network dynamics $\mathcal{N}_{0}(s)$ using Network-Simplification Level 2 in 20, without imposing additional conditions on the controller of GFM 3 (see “no cond.” in Fig. 11). A small-signal load disturbance at node 1 reveals that the closed-loop system remains stable (Fig. 12), confirming the validity of Corollary 1. Likewise, when modeling $\mathcal{N}_{0}(s)$ with Network-Simplification Level 1 in 19 and ensuring that the controller of GFM 3 satisfies the stability conditions in 27 (see “cond. L1” in Fig. 11), we again observe closed-loop stability (Fig. 13). This further confirms Corollary 1. Our main result, Theorem 2, is validated in Fig. 14, where we model $\mathcal{N}_{0}(s)$ dynamically as in 18, and equip GFM 3 with a controller that meets the stability conditions in 25 and 26 (see “cond. DYN” in Fig. 11). Stability is immediately evident. In contrast, using a GFM 3 controller that satisfies only 27 (“cond. L1”) or no conditions at all (“no cond.”) leads to instability in the dynamic network model. This underscores the importance of accurate network modeling in control design, which is overlooked in the overly optimistic stability assessment of [8, 10, 11].
(a) Stability conditions in 25.
(b) Stability conditions in 26.
Figure 11: 2D plot of the stability conditions for the three-bus system with $\rho=0.05$. The stars indicate the feasibility of the GFM 3 control parameters.
Figure 12: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 20, where the controller of GFM 3 does not satisfy any of the conditions 25, 26 and 27 (no cond.).
Figure 13: Stable system response of the block diagram in Fig. 4 with three GFM devices and the simplified network model 19, where the GFM 3 controller satisfies the conditions in 27 (cond. L1).